40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7274     Entry time: Fri Aug 24 21:00:40 2012
Author: Koji 
Type: Update 
Category: LSC 
Subject: X end green investigation 

I checked and fixed the X end green situation. Now the X green beam is locked with TEM00.

There are various reasons it did not lock nicely.

  • The IR beam axis was changed by Yoichi and Rana (ELOG #7169). So the green axis also had to be changed.
  • The end green optics is really "BS". Anytime I see it, I feel disgusted. Because of 3D steering mirrors, cross couplings
    between yaw and pitch are big. This makes the alignment hard.
  • Even with acceptable alignment, the lock was only momentarily. I found the slow control was on. This pushed the frequency
    too much and made the lock unstable.
  • The slow control screen was broken as Jamie changed the model names but did not fix the slow screens.
    • Jamie saids (ELOG #7011): Fix the c1sc{x,y}/master/C1SC{X,Y}_GC{X,Y}_SLOW.adl screens. 
      I need to figure out a more consistent place for those screens.

Now some action items are left:

- IR TRX is not aligned.
- X end green needs precise alignment.
- PSL GR TRX is not aligned.

These will be checked on Sunday.

- End green setup is horrible. => Manasa and I should work on this together.

ELOG V3.1.3-